OpenStudy (anonymous):

how to calculate integral without limits from scientific calculator? If anyone knows ,plz help...its urgent

OpenStudy (anonymous):

use geometric shapes... I am not sure what you are asking for-:(

OpenStudy (anonymous):

are you dealing with a particular problem?

OpenStudy (anonymous):

I want to calculate integration from scientific calculator \[\int\limits\limits x+3x \]

OpenStudy (anonymous):

sure, \[\int\limits_{ }^{ }x+3x~dx\]you mean

OpenStudy (mrnood):

x+3x = 4x

OpenStudy (anonymous):

i can calculate integral with limits on calculator bt dont know how to calculate without limits

OpenStudy (anonymous):

Mr is right, add them, and take the constant out

OpenStudy (anonymous):

\[\int\limits_{ }^{ }x+3x~dx=\int\limits_{ }^{ }4x dx = 4 \times \int\limits_{ }^{ }x ~dx=?\]

OpenStudy (anonymous):

I have to calculte it using calculator.....not by solving. I have randomly written the expression, Actually i want to know the method of solving integration on scientific calculator

OpenStudy (anonymous):

\[Power~rule:~ \int\limits_{ }^{ } x^n~~dx=~~\frac{1}{n+1}x^{n+1}+C\]

OpenStudy (anonymous):

I see, idk how to use the calculator for that. (Nor do I see any purpose in doing that)

OpenStudy (anonymous):

it should be a definite integral, shouldn't?

OpenStudy (anonymous):

we have to solve complex integration questions, which take huge time to solve manually

OpenStudy (anonymous):

but your problem is just an indefinite integral, and you have to use a calculator, or is it a definite integral?

OpenStudy (anonymous):

indefinite

OpenStudy (anonymous):

idk how to sorry-:(

OpenStudy (anonymous):

its ok....n.p

OpenStudy (freckles):

can you post your integral?

OpenStudy (anonymous):

I just want to know the basic method of solving any integral(without limits) on scientific calculator

OpenStudy (anonymous):

freckles it is, as he said.\[\int\limits_{ }^{ }(x+3x)~dx\]

OpenStudy (freckles):

oh i thought it was a little more complex than that

OpenStudy (freckles):

because he said it was really complex

OpenStudy (anonymous):

i have typed this expression randomly to know the method......In exams , several logarthmic and exponential funtions are combined in an expression.......which takes huge time to solve manually

OpenStudy (freckles):

it sometimes depends on the calculator i have only ever known handheld calculators to approximate definite integrals

OpenStudy (anonymous):

I also know how to calculate definite integrals........bt I have to calculate indefinite integrals

OpenStudy (freckles):

have you tried looking up your model of calculator and seeing if it can actually perform this task you want it to or if you want me to look it up i need to know the model of calculator

OpenStudy (anonymous):

fx 991 ES plus fx 991 MS

OpenStudy (freckles):

if i didn't read this wrong the answer is no

OpenStudy (anonymous):

For what it's worth, I have yet to hear of a scientific handheld calculator that can take antiderivatives. As far as I know, most calculators only have a definite-integral function.

OpenStudy (mrnood):

if it is an indefinite integral then the answer is an equation, NOT a value. I can't see how a calculator could do this for any except the most trival equations.

OpenStudy (anonymous):

yes i need the answer as an equation

OpenStudy (freckles):

Wolfram can handle lots of integrals.

OpenStudy (mrnood):

for this particular integral ? or as a general rule? I think you are expecting too much from a calculator - integration can rapidly become a complex process - you need to analyse the method required and the rules and the constant of integration.......

OpenStudy (anonymous):

i need to know general method

OpenStudy (freckles):

@1234sd I looked up your model and a bunch people said you couldn't do it with that model.

OpenStudy (zzr0ck3r):

for the integral \(\int_a^bf(x) \ dx\) On a TI-89 `int(f(x), x, a, b)` for \(\int f(x) \ dx\) use `int(f(x),x)` int is the little integral sign above the 7

OpenStudy (zzr0ck3r):

it will not retricethe constant, so you need to do that.... i.e. for \(\int 2x \ dx\) it will give the answer as \(x^2\) you need to add the constant so \(x^2+c\)

OpenStudy (zzr0ck3r):

@SithsAndGiggles ti 89 will most definitely handle indefinite integrals.

OpenStudy (anonymous):

I would have expected it as a feature on the NSpire model, but not the TI-89. Good to know! I guess I hadn't considered graphing calculators as "scientific" for some reason.